# Randomization Tests for Conditional Group Symmetry

### v1

@@ B.2 Proofs for Section 4
Let $(X,Y)\sim P_{X,Y}$. Assume that $P_{Y|X}$ is
$\mathbf{G}$-equivariant. To show Property E1, let $P_{[\mathbf{X}]}=\rho_{*}P_{X}$, so that for any integrable $f\colon[\mathbf{X}]\times\mathbf{Y}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$,
$\int$ $f(\rho(x),\tilde{y},x,y)P(dx,d\tilde{y},dy)$
$=\int f([x],\tilde{y},x^{\prime},y)P_{Y|X}(dy\mid\tilde{y},x^{\prime})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{X|[X]}(dx^{\prime}\mid[x])P_{[X]}(d[x])$
$=\int f([x],\tilde{y},\tilde{\tau}[x],\tilde{\tau}\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x^{\prime})P_{X|[X]}(dx^{\prime}\mid[x])P_{[X]}(d[x])$
$=\int f([x],\tilde{y},\gamma[x],\gamma\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{\Gamma|[X]}(d\gamma\mid[x])P_{[X]}(d[x])$
$=\int f(\rho(x),\tilde{y},\gamma\rho(x),\gamma\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid\rho(x))P_{\Gamma|[X]}(d\gamma\mid\rho(x))P_{X}(dx)\;,$
To show that Property E1 implies that $P_{Y|X}$ is $\mathbf{G}$-equivariant, observe that $(\rho(X),\tilde{Y},X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\rho(X),\tilde{Y},\Gamma\rho(X),\Gamma\tilde{Y})$ implies that $(X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\Gamma\rho(X),\Gamma\tilde{Y})$, where $\Gamma,\tilde{Y}$ are sampled as in (14). By Lemma B.4, that implies that $P_{Y|X}$ is $\mathbf{G}$-equivariant.
Proof of Corollary 4.3.
Property E4Property E6 follow as special cases of the corresponding properties in Proposition 4.2, marginalizing over $X$. ∎
Proof of Corollary 4.4.
Properties E7 and E8 follow from the fact that
$P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|[x])=\lambda({\,\vbox{\hbox{\tiny$\bullet$}}\,})$ for each $[x]$ if and only if $P_{X}$ is $\mathbf{G}$-invariant. ∎
B.3 Proofs for Section 5
Proof of Proposition 5.1.
By Proposition 4.2, each of the randomization procedures generate conditionally i.i.d. samples, conditioned on $S_{1:n}$. By construction, the randomized samples have the same distribution as the observed samples under the null hypothesis. Then (17) follows from Lemma B.3.
@@ B.3 Proofs for Section 5
Proof of Proposition 5.2.
@@
B.4 Proofs for Section 6
B.4.1 Proof of Theorem 6.1
@@ B.4 Proofs for Section 6
$\mathrm{MMD}_{\sigma}(P,Q)^{2}=\left(\sup_{f\in\mathcal{H}|_{1}}\left|\int f(x)P(dx)-\int f(x)Q(dx)\right|\right)^{2}$
$=\mathbb{E}[k(X,\tilde{X})]+\mathbb{E}[k(X^{\prime},\tilde{X}^{\prime})]-2\mathbb{E}[k(X,X^{\prime})]\;,\quad X,\tilde{X}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P\;,X^{\prime},\tilde{X^{\prime}}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}Q\;,$
@@
$\mathrm{MMD}_{\sigma}\textbf{}(\hat{P},\hat{Q})^{2}=\left(\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-f(X^{\prime}_{i}))\right|\right)^{2}\;,\quad X_{i}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P\;,\ X^{\prime}_{i}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}Q\;.$
@@
  • $\int_{\mathbf{X}}\theta_{i}(f,x)^{2}P_{i}(dx)\leq\sigma^{2}_{i}<\infty$; and
@@
$P\left\{w\in\mathbf{W}:H(w)-\mathbb{E}[H(W)]\geq\epsilon\right\}\leq e^{-\frac{3}{2}\frac{\epsilon^{2}}{3a+\epsilon}}\;,$ (25)
with $a=2M\mathbb{E}[H(W)]+\sum_{i=1}^{n}\sigma^{2}_{i}$. Written differently, for all $\eta>0$,
$P\left\{w\in\mathbf{W}:H(w)\geq\mathbb{E}[H(W)]+\sqrt{2a\eta}+2\eta M/3\right\}\leq e^{-\eta}\;.$ (26)
@@
$T(\hat{P}_{1:n},\hat{Q}_{1:n})=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}b_{i}f(Z_{i})\right|=\left(\frac{1}{n^{2}}\sum_{i=1,j=1}^{2n}b_{i}b_{j}k(Z_{i},Z_{j})\right)^{1/2}\;.$ (27)
@@
$T(P_{1:n},Q_{1:n})=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}b_{i}\mathbb{E}[f(Z_{i})]\right|=\left(\frac{1}{n^{2}}\sum_{i=1,j=1}^{2n}b_{i}b_{j}\mathbb{E}[k(Z_{i},Z_{j})]\right)^{1/2}\;,$ (28)
which is equal to 00 if $P_{i}=Q_{i}$ for each $i\in[n]$. Finally, observe that with
$\theta_{i}(f,x)=(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}\frac{1}{n}\bigg{(}f(x)-\int f(x^{\prime})P_{i}(dx^{\prime})\bigg{)}\;,\quad f\in\mathcal{H}|_{1}\;,$ (29)
the conditions of Lemma B.5 are satisfied, with $\sigma_{i}^{2}\leq\nu^{2}/n^{2}$ and $M\leq 4\nu/n$.
@@
$|T(\hat{P}_{1:n},\hat{Q}_{1:n})-T(P_{1:n},Q_{1:n})|$
$\qquad\leq\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}(f(Z_{i})-\mathbb{E}[f(Z_{i})])\right|$
$\qquad=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}\theta_{i}(f,Z_{i})\right|\;.$
We will make repeated use of the following upper bound on the expectation of $H(W)$, a proof of which can be found in, for example, (Sriperumbudur et al., 2012, Appendix D).
Lemma B.6.
@@
$\mathbb{E}[H(W)]\leq 2\nu\sqrt{\frac{2}{n}}\;.$
The factor of $2$ comes from a symmetrization argument (Sriperumbudur et al., 2012, Appendix B.2). The rest of the upper bound is a bound on the Rademacher complexity of the RKHS.
@@
Let $P_{i}$ be the conditional distribution of $Z_{i}$ conditioned on $S_{i}=s_{i}$, and let $Q_{i}=P_{i}$ be the same for $Z^{\prime}_{i}$. In particular, with $\hat{\Delta}_{n}=|T(\hat{P},\hat{Q}^{(0)})|$ and $\bar{H}_{n}=\mathbb{E}[H(W)]$, if $\hat{\Delta}_{n}>\bar{H}_{n}$ then
@@
$p(\hat{\Delta}_{n})\coloneq\;$ :- $\mathbb{P}\left(\left|T(\hat{P}^{(b)},\hat{Q}^{(b)})\right|\geq\hat{\Delta}_{n}\mid S_{1:n},\hat{\Delta}_{n}\right)$
$\leq\exp\left(-\frac{3}{2}\frac{(\hat{\Delta}_{n}-\bar{H}_{n})^{2}}{3a+\hat{\Delta}_{n}-\bar{H}_{n}}\right)$
$\leq\exp\left(-\frac{3}{2}(\hat{\Delta}_{n}-\bar{H}_{n})+\frac{9}{2}a\right)$
$\leq\exp\left(-\frac{3}{2}\hat{\Delta}_{n}+\frac{9}{2n}\nu^{2}+\frac{16\nu^{2}\sqrt{2}}{\sqrt{n^{3}}}+3\nu\sqrt{\frac{2}{n}}\right)=:\hat{p}(\hat{\Delta}_{n})\;,$
where the first inequality follows from Lemma B.5, the second inequality follows from $e^{-1/(1+x)}\leq e^{-(1-x)}$, $x\geq 0$, and the third inequality follows from $\sigma^{2}_{i}\leq\nu^{2}/n^{2}$ and Lemma B.6.
Now, conditioned on $\hat{\Delta}_{n}$ and $S_{1:n}$, $N_{B}\coloneq p_{B}(B+1)-1$ is a random variable with $\textrm{Binomial}(B,p(\hat{\Delta}_{n}))$ distribution, so that
$\mathbb{P}(p_{B}\leq\alpha$ $\mid S_{1:n},\hat{\Delta}_{n})=\mathbb{P}(N_{B}\leq\alpha(B+1)-1\mid S_{1:n},\hat{\Delta}_{n})$

### v2

@@ C.3 Proofs for Section 4
By Proposition C.1, each of the randomization procedures generates conditionally i.i.d. samples, conditioned on $S_{1:n}$. By construction, the randomized samples have the same distribution as the observed samples under the null hypothesis. Then (13) follows from Lemma B.4.
@@
Proposition C.5.
Assume $(X,Y)\sim P_{X,Y}$, with $P_{Y|X}$ $\mathbf{G}$-equivariant. Let $T$ be a continuous $\mathbb{R}$-valued function, such that the distribution of $T((X,Y)_{1:n})$ is atomless. Suppose that $(X,Y)_{i,m}^{(b)}$ are sampled conditionally on $S(X,Y)_{i,m}^{(b)}=S(X,Y)_{i}$, and used as the randomization sets to generate CRT statistics as in (11), denoted $p_{B,m}$. If, for each $i\leq n$, conditioned on $S(X,Y)_{i}$, $(X,Y)_{i,m}^{(b)}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}(X,Y)_{i}$ as $m\to\infty$, and $\mathbb{P}\{T((X,Y)_{1:n}^{(b)})=T((X,Y)_{1:n}^{(b^{\prime})})\}=0$ for $b\neq b^{\prime}$, then $p_{B,m}\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{\to}}p_{B}$. Therefore, for fixed $B\in\mathbb{N}$ and $\alpha\in[0,1]$,
$\lim_{m\to\infty}\mathbb{P}\left(p_{B,m}\leq\alpha\mid S_{1:n}\right)\leq\alpha\;,\quad P\text{-a.s.}$
Proof.
@@
C.4 Additional results and proofs for Section 5
C.4.1 Proof of Theorem 5.1
@@ C.4 Additional results and proofs for Section 5
$\mathrm{MMD}_{\sigma}^{2}$ $(P,Q)=\left(\sup_{f\in\mathcal{H}|_{1}}\left|\int f(x)P(dx)-\int f(x)Q(dx)\right|\right)^{2}$
$=\mathbb{E}[k_{\sigma}(X,\tilde{X})]+\mathbb{E}[k_{\sigma}(X^{\prime},\tilde{X}^{\prime})]-2\mathbb{E}[k_{\sigma}(X,X^{\prime})]\;,\quad X,\tilde{X}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P\;,X^{\prime},\tilde{X^{\prime}}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}Q\;,$
@@
$\mathrm{MMD}_{\sigma}^{2}(X_{1:n},X_{1:n}^{\prime})=\left(\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-f(X^{\prime}_{i}))\right|\right)^{2}\;,\quad X_{i}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}P\;,\ X^{\prime}_{i}\stackrel{{\scriptstyle\text{\rm\tiny iid}}}{{\sim}}Q\;.$
@@
  • $\int_{\mathbf{X}}\theta_{i}(f,x)^{2}P_{i}(dx)\leq\varphi^{2}_{i}<\infty$; and
@@
$P\left(w\in\mathbf{W}:H(w)-\mathbb{E}[H(W)]\geq\epsilon\right)\leq e^{-\frac{3}{2}\frac{\epsilon^{2}}{3a+\epsilon}}\;,$
with $a=2M\mathbb{E}[H(W)]+\sum_{i=1}^{n}\varphi^{2}_{i}$. Written differently, for all $\eta>0$,
$P\left(w\in\mathbf{W}:H(w)\geq\mathbb{E}[H(W)]+\sqrt{2a\eta}+2\eta M/3\right)\leq e^{-\eta}\;.$
@@
$T(Z_{1:n},Z_{(n+1):2n})=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}b_{i}f(Z_{i})\right|=\left(\frac{1}{n^{2}}\sum_{i=1,j=1}^{2n}b_{i}b_{j}k_{\sigma}(Z_{i},Z_{j})\right)^{1/2}\;.$
@@
$T(P_{1:n},Q_{1:n})=\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}b_{i}\mathbb{E}[f(Z_{i})]\right|=\left(\frac{1}{n^{2}}\sum_{i=1,j=1}^{2n}b_{i}b_{j}\mathbb{E}[k_{\sigma}(Z_{i},Z_{j})]\right)^{1/2}\;,$
which is equal to 0 if $P_{i}=Q_{i}$ for each $i\in[n]$. Finally, observe that with
$\theta_{i}(f,z)=(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}\frac{1}{n}\bigg(f(z)-\int f(z^{\prime})P_{i}(dz^{\prime})\bigg)\;,\quad f\in\mathcal{H}|_{1}\;,$ (20)
the conditions of Lemma C.6 are satisfied, with $\varphi_{i}^{2}\leq\kappa^{2}/n^{2}$ and $M\leq 4\kappa/n$.
@@
$\left|T(Z_{1:n},Z_{(n+1):2n})-T(P_{1:n},Q_{1:n})\right|$
$\quad=\frac{1}{n}\left|\sup_{f\in\mathcal{H}\mid_{1}}\left|\sum_{i=1}^{2n}(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}f(Z_{i})\right|-\sup_{f\in\mathcal{H}\mid_{1}}\left|\sum_{i=1}^{2n}(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}\mathbb{E}[f(Z_{i})]\right|\right|$
$\quad\leq\sup_{f\in\mathcal{H}\mid_{1}}\frac{1}{n}\left|\sum_{i=1}^{2n}(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}\left(f(Z_{i})-\mathbb{E}[f(Z_{i})]\right)\right|$ (21)
$\quad=\sup_{f\in\mathcal{H}\mid_{1}}\left|\sum_{i=1}^{2n}\theta_{i}(f,Z_{i})\right|\;.$
We will make repeated use of the following upper bound on the expectation of $H(W)$, the proof of which is standard, for example, (Sriperumbudur et al., 2012, Appendix D).
Lemma C.7.
@@
$\mathbb{E}[H(W)]\leq 2\kappa\sqrt{\frac{2}{n}}\;.$
Proof of Lemma C.7.
For independent $Z_{i},Z_{i}^{\prime}\sim P_{i}$ and $Z_{n+i},Z_{n+i}^{\prime}\sim Q_{i}$, and independent Rademacher random variables $R_{i}$, $i\in[n]$,
$\mathbb{E}[H(W)]$ $=\mathbb{E}\left[\sup_{f\in\mathcal{H}|_{1}}\left|\sum_{i=1}^{2n}\theta_{i}(f,Z_{i})\right|\right]$
$=\mathbb{E}\left[\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}\left(f(Z_{i})-\mathbb{E}[f(Z_{i})]\right)\right|\right]$
$\leq\mathbb{E}\left[\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}(-1)^{\mathbbm{1}\{n+1\leq i\leq 2n\}}\left(f(Z_{i})-f(Z_{i}^{\prime})\right)\right|\right]$
$=\mathbb{E}\left[\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}R_{i}\left(f(Z_{i})-f(Z_{i}^{\prime})\right)\right|\right]$
$\leq 2\mathbb{E}\left[\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}R_{i}f(Z_{i})\right|\right]\;,$
where the first inequality follows from Jensen’s inequality. Using RKHS identities and the boundedness of the kernel (see Appendix D, Sriperumbudur et al., 2012) yields
$\mathbb{E}\left[\sup_{f\in\mathcal{H}|_{1}}\left|\frac{1}{n}\sum_{i=1}^{2n}R_{i}f(Z_{i})\right|\right]\leq\kappa\sqrt{\frac{2n}{n^{2}}}=\kappa\sqrt{\frac{2}{n}}\;.$
@@
The proof proceeds in two stages. We first derive a high-probability upper bound (over observations $Z_{1:n}$ and randomization samples $Z_{1:n}^{(0)}$) for the conditional probability that a randomized test statistic is greater than the observed test statistic, conditional on $\mathcal{A}_{1:n}\coloneqq(S_{1:n},Z_{1:n},Z_{1:n}^{(0)})$. We then use the probability upper bound to lower bound the test power in terms of a binomial cumulative distribution function.
[…]
$p(\hat{\Delta}_{n})$ $\coloneqq\mathbb{P}\left(\left.\left|T(Z^{(b)}_{1:n},Z^{\prime(b)}_{1:n})\right|\geq\hat{\Delta}_{n}\;\right|\;\mathcal{A}_{1:n}\right)$
$=\mathbb{P}\left(\left.\left|T(Z^{(b)}_{1:n},Z^{\prime(b)}_{1:n})\right|\geq\hat{\Delta}_{n}\;\right|\;\mathcal{A}_{1:n},\hat{\Delta}_{n}>\xi_{n}\right)\mathbbm{1}\left\{\hat{\Delta}_{n}>\xi_{n}\right\}$
@@
$\quad+\mathbb{P}\left(\left.\left|T(Z^{(b)}_{1:n},Z^{\prime(b)}_{1:n})\right|\geq\hat{\Delta}_{n}\;\right|\;\mathcal{A}_{1:n},\hat{\Delta}_{n}\leq\xi_{n}\right)\mathbbm{1}\left\{\hat{\Delta}_{n}\leq\xi_{n}\right\}$
$\leq\min\left\{1,\mathbb{P}\left(\left.\left|T(Z^{(b)}_{1:n},Z^{\prime(b)}_{1:n})\right|\geq\hat{\Delta}_{n}\;\right|\;\mathcal{A}_{1:n},\hat{\Delta}_{n}>\xi_{n}\right)+\mathbbm{1}\left\{\hat{\Delta}_{n}\leq\xi_{n}\right\}\right\}\;.$
Focusing on the first term in the upper bound, we have
$p^{*}$ $\coloneqq\mathbb{P}\left(\left.\left|T(Z^{(b)}_{1:n},Z^{\prime(b)}_{1:n})\right|\geq\hat{\Delta}_{n}\;\right|\;\mathcal{A}_{1:n},\hat{\Delta}_{n}>\xi_{n}\right)$
$=\mathbb{P}\left(\left.\left|T(Z^{(b)}_{1:n},Z^{\prime(b)}_{1:n})\right|-\xi_{n}\geq\hat{\Delta}_{n}-\xi_{n}\;\right|\;\mathcal{A}_{1:n},\hat{\Delta}_{n}>\xi_{n}\right)$
$\leq\exp\left(-\frac{3}{2}\frac{(\hat{\Delta}_{n}-\xi_{n})^{2}}{3a+\hat{\Delta}_{n}-\xi_{n}}\right)$
$=\exp\left(-\frac{3}{2}\left(\hat{\Delta}_{n}-\xi_{n}-3a+\frac{9a^{2}}{3a+\hat{\Delta}_{n}-\xi_{n}}\right)\right)$
$\leq\min\left\{1,\exp\left(-\frac{3}{2}\hat{\Delta}_{n}+\frac{3}{2}\xi_{n}+\frac{9}{2}a\right)\right\}$ (22)
$\leq\min\left\{1,\exp\left(-\frac{3}{2}\hat{\Delta}_{n}+3\kappa\sqrt{\frac{2}{n}}+72\kappa^{2}\sqrt{\frac{2}{n^{3}}}+\frac{9}{n}\kappa^{2}\right)\right\}\;.$
The first inequality follows from Lemma C.6 and the fact that the conditional distributions of $Z_{1:n}^{(b)}$ and $Z_{1:n}^{\prime(b)}$ are the same. The third inequality follows from the bound $a\leq\frac{2\kappa^{2}}{n}\left(1+8\sqrt{\frac{2}{n}}\right)$, which comes from Lemma C.7 and the bounds on $\varphi_{i}^{2}$ and $M$.
From Lemma C.6 conditioned on $S_{1:n}$, for $\eta>0$, we have with probability over $Z_{1:n}$ and $Z_{1:n}^{(0)}$,
$1-e^{-\eta}$ $\leq\mathbb{P}\left(\left.H(Z_{1:2n})\leq\xi_{n}+\sqrt{2a\eta}+\frac{2M\eta}{3}\;\right|\;S_{1:n}\right)$
$\leq\mathbb{P}\left(\left.\left|T(P_{1:n},P^{0}_{1:n})-T(Z_{1:n},Z^{(0)}_{1:n})\right|\leq\xi_{n}+\sqrt{2a\eta}+\frac{2M\eta}{3}\;\right|\;S_{1:n}\right)$
$\leq\mathbb{P}\left(\left.T(P_{1:n},P^{0}_{1:n})-\left|T(Z_{1:n},Z^{(0)}_{1:n})\right|\leq\xi_{n}+\sqrt{2a\eta}+\frac{2M\eta}{3}\;\right|\;S_{1:n}\right)$
$\leq\mathbb{P}\left(\left.\Delta_{n}-\hat{\Delta}_{n}\leq 2\kappa\sqrt{\frac{2}{n}}+\sqrt{2\eta\left(\frac{2\kappa^{2}}{n}\left(1+8\sqrt{\frac{2}{n}}\right)\right)}+\frac{8\kappa\eta}{3n}\;\right|\;S_{1:n}\right)$
$=\mathbb{P}\left(\left.\Delta_{n}-2\kappa\left(\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\leq\hat{\Delta}_{n}\;\right|\;S_{1:n}\right)\;,$ (23)
where the second inequality follows from (21), the third inequality from the fact $T(P_{1:n},P^{0}_{1:n})\geq 0$, and the fourth inequality from applying the previous bounds on $\xi_{n}$, $a$, and $M$. Applying the probabilistic bound in (23) to the upper bound on $p^{*}$, we find that with probability at least $1-e^{-\eta}$ over $Z_{1:n}$ and $Z_{1:n}^{(0)}$,
[…]
+ $\qquad\qquad\left.\left.+\;3\kappa\sqrt{\frac{2}{n}}+72\kappa^{2}\sqrt{\frac{2}{n^{3}}}+\frac{9}{n}\kappa^{2}\right)\right\}$
$=\min\left\{1,\exp\left(-\frac{3}{2}\Delta_{n}+C_{n}(\eta)\right)\right\}$
$=:p_{\eta}(\Delta_{n})\;,$
where $C_{n}(\eta)$ is defined as in (14). Therefore,
$p(\hat{\Delta}_{n})$ $\leq\min\left\{1,p_{\eta}(\Delta_{n})+\mathbbm{1}\left\{\hat{\Delta}_{n}\leq\xi_{n}\right\}\right\}\;.$
By (23) and Lemma C.7, with the same probability,
$\mathbbm{1}\left\{\hat{\Delta}_{n}\leq\xi_{n}\right\}$ $\leq\mathbbm{1}\left\{\Delta_{n}-2\kappa\left(\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\leq 2\kappa\sqrt{\frac{2}{n}}\right\}$
$=\mathbbm{1}\left\{\Delta_{n}\leq 2\kappa\left(2\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\right\}\;,$
which equals zero under assumption (15). Hence, $p(\hat{\Delta}_{n})\leq p_{\eta}(\Delta_{n})$ with probability at least $1-e^{-\eta}$.
[…]
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})$ $=\mathbb{P}(N_{B}\leq\alpha(B+1)-1\mid S_{1:n})$